OpenStudy (anonymous):

has endpoints at -2 and 12. What is the coordinate of its midpoint?

OpenStudy (anonymous):



OpenStudy (anonymous):

differnce bet ween points so in this case -2-12 then divide that by 2 and you will hav ea mid point coodrinate but you need an x and y component you have youonly given on coordinate

OpenStudy (amistre64):

-2 to 12 if we add 2 to each end -2 to 12 +2 +2 --------- 0 to 14 what is half of 14?

OpenStudy (anonymous):

7?

OpenStudy (amistre64):

correct so ... lets put that in and readjust it back 0 to 7 to 14 -2 -2 -2 ----------- -2 to 5 to 12

OpenStudy (amistre64):

we get the same thing by using the midpoint formula: \[\frac{first+last}{2}\]\[\frac{-2+12}{2}=10/2 = 5\]